# InvAgent: A Large Language Model based Multi-Agent System for Inventory Management in Supply Chains

### v1

@@ 4.2 Baselines
Base-Stock Policy. The base-stock policy Lee et al. (1997); Oroojlooyjadid et al. (2022) is an inventory management strategy where each stage in the supply chain maintains a constant inventory level, or base-stock level. Here, the desired inventory level is set equal to the production capacity:
@@
More heuristic baselines are discussed in Appendix B.
@@
Independent Proximal Policy Optimization (IPPO) with Parameter Sharing. Proximal policy optimization (PPO) Schulman et al. (2017) updates policies by iteratively sampling data through interaction with the environment and optimizing a clipped surrogate objective function using multiple epochs of stochastic gradient ascent. Independent PPO (IPPO) De Witt et al. (2020) is a RL approach where each agent is independently trained with the PPO algorithm. Here we employ the parameter sharing by using the same policy parameters for all agents to improve learning efficiency and coordination.
@@
Multi-Agent Proximal Policy Optimization (MAPPO). Multi-agent PPO (MAPPO) Yu et al. (2022) is an extension of PPO algorithm designed for multi-agent environments. It utilizes a centralized value function that takes into account global information from all agents, improving variance reduction and stability during training.
4.3 Experiment Settings
The performance of our model InvAgent is evaluated using the total reward from all stages all periods during one simulation (episode), and the reported numbers are averaged over 5 episodes for each experiment to reduce the uncertainty. We utilize the Python packages such as AutoGen Wu et al. (2023), Gymnasium Towers et al. (2023), and RLlib Liang et al. (2018), and also LLMs including GPT-4, GPT-4O, and GPT-4-Turbo Achiam et al. (2023). For the constant demand scenario, we change the last part of the prompt in Figure 4 to "([0], [4], or [8] only)" to boost the InvAgent performance.
@@ 4.3 Experiment Settings
The performance of baseline models is evaluated based on the episode reward averaged over 100 episodes. For reinforcement learning (RL), we explore various hyper-parameter settings, including the numbers of hidden units ([128, 128] and [256, 256]), activation function (ReLU), learning rate (1e-4, 5e-4, 1e-3), training batch size (500, 1000, 2000), stochastic gradient descent (SGD) minibatch size (32, 64, 128), number of SGD iterations (5, 10, 20), number of training iterations (500, 800, 1000, 1500), and discount factor (1.0). We randomly select 20 combinations of these hyper-parameters during one experiment and keep the best one of them. The final hyper-parameters used for all scenarios are presented in Appendix C. The RL experiments are conducted on a NVIDIA A10 GPU.
4.4 Experiment Results
-2cm-2cm Model Constant Variable Larger Seasonal Normal Base-Stock -296.00 (0.00) -523.69 (49.15) -392.21 (111.79) -274.29 (40.75) -322.44 (99.59) Tracking-Demand -360.00 (0.00) -412.41 (41.76) -265.07 (99.67) -421.90 (55.18) -232.20 (75.45) IPPO -132.17 (40.17) -389.55 (40.28) -202.39 (92.96) -126.73 (183.63) -102.90 (64.68) MAPPO -129.81 (16.02) -391.53 (34.09) -106.79 (109.86) -99.39 (126.09) -41.98 (75.22) InvAgent (w/o strategy) -156.00 (0.00) -336.60 (43.24) -350.20 (149.57) -488.00 (114.82) -172.60 (104.70) InvAgent (w/ strategy) -200.00 (0.00) -377.60 (53.50) -357.60 (50.04) -420.60 (225.42) -192.40 (98.51)
When compared to heuristic baselines, InvAgent shows notable advantages. Unlike the base-stock policy and tracking-demand policy, which rely on fixed or historical data and often struggle with fluctuating demands, InvAgent dynamically adapts to real-time conditions, minimizing inventory costs and avoiding stockouts. InvAgent performs particularly well in the variable demand scenario, indicating its effectiveness in managing unpredictable demand patterns. These results demonstrate that InvAgent can achieve lower costs and better adaptability than the heuristic baselines.
In comparison to RL models, InvAgent offers distinct strengths despite certain performance limitations. RL models like MAPPO and IPPO achieve higher metrics in several scenarios due to extensive training but come with increased complexity and potential instability. RL training often requires significant computational resources and time, with risks of overfitting. In contrast, InvAgent provides key advantages in explainability and ease of implementation, making reasonable decisions without prior training. Although InvAgent does not always outperform RL models, its stability, simplicity, and interpretability make it a valuable alternative for dynamic inventory management.
The comparison of the InvAgent model with and without strategy shows that the inherent strategy of the LLM is generally sufficient for many simpler scenarios, such as constant and variable demands. However, in more complex scenarios like seasonal demands, the addition of a human-crafted strategy enhances the LLM’s decision-making. This demonstrates that while the model’s built-in capabilities are adequate for straightforward situations, incorporating explicit strategies proves beneficial for managing more complex and patterned demand scenarios, thereby improving performance and adaptability.
4.5 Ablation Studies
We conduct ablation studies to evaluate the impact of various prompt components in InvAgent under the variable demand scenario, as detailed in Table 4. This scenario introduces randomness with demand varying uniformly between 0 and 4 units per period over 12 periods. The results demonstrate that the contributions of different components to the model’s performance vary significantly. The prompt without the strategy component is the best one and other models are compared with it.
-2cm-2cm Model Demand Downstream Strategy CoT History Reward % GPT-4 ✓ ✓ ✗ ✓ ✓ -336.60 (43.24) 0.00% GPT-4 ✓ ✓ ✓ ✓ ✓ -377.60 (53.50) -12.18% GPT-4 ✗ ✓ ✓ ✓ ✓ -349.40 (29.43) -3.80% GPT-4 ✓ ✗ ✓ ✓ ✓ -419.00 (35.91) -24.48% GPT-4 ✗ ✗ ✓ ✓ ✓ -379.40 (40.03) -12.72% GPT-4 ✓ ✓ ✗ ✓ ✗ -339.20 (46.63) -0.77% GPT-4 ✓ ✓ ✓ ✗ ✓ -369.80 (36.83) -9.86% GPT-4 ✓ ✓ ✓ ✓ ✗ -387.40 (11.09) -15.09% GPT-4O ✓ ✓ ✓ ✓ ✓ -405.00 (35.14) -20.32% GPT-4-Turbo ✓ ✓ ✓ ✓ ✓ -636.40 (195.26) -89.07%
Overall, the ablation study confirms the robustness and adaptability of our model in dynamic supply chain environments. Components such as the demand description and downstream order are particularly essential for optimizing performance under variable demands. By keeping agent histories, all previous messages in the entire episode (simulation) are retained, allowing stage agents to use the entire chat history for context in their decision-making. Additionally, structured reasoning through Chain-of-Thought (CoT) also play crucial roles. These findings emphasize the importance of each component in achieving effective and efficient inventory management, guiding future improvements and applications in more complex scenarios.
5 Conclusion
In this study, we demonstrate the effectiveness of using large language models (LLMs) as autonomous agents for multi-agent inventory management in supply chain optimization. Our novel model, InvAgent, leverages the zero-shot learning capabilities of LLMs, enabling them to make adaptive and informed decisions without prior training. The integration of structured reasoning through the Chain-of-Thought (CoT) methodology further enhances the explainability and transparency of the model, making it more reliable and easier to trust compared to traditional heuristic or reinforcement learning models. The experimental results show that our model performs competitively, achieving lower costs and better adaptability compared with heuristic polices across various demand scenarios. This highlights the potential of LLMs to notably improve supply chain management by reducing inventory costs and minimizing stockouts.
In the future, we will fine-tune our model using reinforcement learning to enhance decision-making capabilities, allowing the LLMs to learn and optimize strategies over iterations. We also plan to use real-world data to evaluate the efficiency of our model and the utility of the agents. For real data with seasonality, decomposing data into level, trend, and seasonality components will be explored to refine predictive accuracy. Combining human-crafted strategies with the LLMs’ inherent capabilities will be another focus to handle varied and unpredictable demand patterns more robustly.
Limitations
Scope of Data. The primary limitation of this study is the dependence on simulated scenarios and synthetic data, which may not fully capture the complexities and variabilities of real-world supply chain environments. This reliance restricts the generalizability of our findings to actual supply chain operations, which may exhibit different behaviors and challenges.
Computational Resources. In our model, we need to call the OpenAI API, which incurs significant costs. Despite this limitation, our approach offers valuable insights into the use of large language models for dynamic inventory management, suggesting the need for further research and refinement.
Ethical Considerations
This study adheres to ethical AI principles, ensuring transparency, fairness, and accountability. We use only synthetic data, avoiding any private or sensitive information. Rigorous standards are maintained in experiment design and result evaluation.
Appendix A Prompt and Response Example
Appendix B Evaluation Results of Heuristic Baseline Variants
@@
-2cm-2cm Desired Inventory Constant Variable Larger Seasonal Normal -208.00 (0.00) -435.69 (49.15) -234.28 (102.81) -207.75 (34.67) -150.67 (101.80) -252.00 (0.00) -479.69 (49.15) -310.74 (109.16) -229.08 (34.66) -226.31 (103.32) -296.00 (0.00) -523.69 (49.15) -392.21 (111.79) -274.29 (40.75) -322.44 (99.59) -364.00 (0.00) -390.17 (44.24) -393.31 (79.05) -525.84 (47.85) -283.39 (61.83) -120.00 (0.00) -395.68 (41.86) -470.55 (76.73) -524.26 (64.68) -351.23 (90.08) -360.00 (0.00) -412.41 (41.76) -265.07 (99.67) -421.90 (55.18) -232.20 (75.45) -252.00 (0.00) -382.77 (48.50) -489.75 (110.96) -610.03 (94.43) -177.54 (70.87) -361.00 (0.00) -397.22 (50.02) -325.81 (98.39) -479.07 (69.47) -218.98 (73.26)
Appendix C Reinforcement Learning Baseline Settings
The hyperparameter settings used for the independent proximal policy optimization (IPPO) with parameter sharing and multi-agent proximal policy optimization (MAPPO) baselines are provided in Table 6 and Table 7 respectively.
-2cm-2cm Hyperparameter Constant Variable Larger Seasonal Normal Numbers of Hidden Unit [128, 128] [256, 256] [128, 128] [128, 128] [128, 128] Activation Function ReLU ReLU ReLU ReLU ReLU Learning Rate 0.0001 0.0001 0.001 0.0005 0.0005 Training Batch Size 1000 1000 2000 2000 1000 SGD Minibatch Size 128 128 128 128 128 Number of SGD Iterations 5 10 5 5 5 Number of Training Iterations 1000 1500 1000 800 500
-2cm-2cm Hyperparameter Constant Variable Larger Seasonal Normal Numbers of Hidden Unit [128, 128] [128, 128] [128, 128] [256, 256] [128, 128] Activation Function ReLU ReLU ReLU ReLU ReLU Learning Rate 0.0001 0.0001 0.001 0.0001 0.0001 Training Batch Size 500 2000 2000 1000 500 SGD Minibatch Size 128 32 32 128 128 Number of SGD Iterations 10 5 10 10 10 Number of Training Iterations 500 500 800 1500 500
Appendix D Case Studies
@@
D.1 Variable Demand Scenario
@@
D.2 Seasonal Demand Scenario

### v2

@@ B.2 More Details about Baselines
Base-Stock Policy. The base-stock policy (Lee, Padmanabhan, and Whang 1997; Oroojlooyjadid et al. 2022) is an inventory management strategy where each stage in the supply chain maintains a constant inventory level, or base-stock level. Here, the desired inventory level is set equal to the production capacity:
@@
More heuristic baselines and their results are discussed in Appendix B.4.
@@
Independent Proximal Policy Optimization (IPPO) with Parameter Sharing. Proximal policy optimization (PPO) (Schulman et al. 2017) updates policies by iteratively sampling data through interaction with the environment and optimizing a clipped surrogate objective function using multiple epochs of stochastic gradient ascent. Independent PPO (IPPO) (De Witt et al. 2020) is a RL approach where each agent is independently trained with the PPO algorithm. Here we employ the parameter sharing by using the same policy parameters for all agents to improve learning efficiency and coordination.
@@
Multi-Agent Proximal Policy Optimization (MAPPO). Multi-agent PPO (MAPPO) (Yu et al. 2022) is an extension of PPO algorithm designed for multi-agent environments. It utilizes a centralized value function that takes into account global information from all agents, improving variance reduction and stability during training.
B.3 Reinforcement Learning Baseline Settings
@@
For reinforcement learning (RL), we explore various hyper-parameter settings, including the numbers of hidden units ([128, 128] and [256, 256]), activation function (ReLU), learning rate (1e-4, 5e-4, 1e-3), training batch size (500, 1000, 2000), stochastic gradient descent (SGD) minibatch size (32, 64, 128), number of SGD iterations (5, 10, 20), number of training iterations (500, 800, 1000, 1500), and discount factor (1.0). We randomly select 20 combinations of these hyperparameters during one experiment and keep the best one of them. The final hyperparameter settings used for the independent proximal policy optimization (IPPO) with parameter sharing and multi-agent proximal policy optimization (MAPPO) baselines are provided in Table 5 and Table 6 respectively.
Hyperparameter | Constant | Variable | Larger | Seasonal | Normal |
Numbers of Hidden Unit | [128, 128] | [256, 256] | [128, 128] | [128, 128] | [128, 128] |
Activation Function | ReLU | ReLU | ReLU | ReLU | ReLU |
Learning Rate | 0.0001 | 0.0001 | 0.001 | 0.0005 | 0.0005 |
Training Batch Size | 1000 | 1000 | 2000 | 2000 | 1000 |
SGD Minibatch Size | 128 | 128 | 128 | 128 | 128 |
Number of SGD Iterations | 5 | 10 | 5 | 5 | 5 |
Number of Training Iterations | 1000 | 1500 | 1000 | 800 | 500 |
Hyperparameter | Constant | Variable | Larger | Seasonal | Normal |
Numbers of Hidden Unit | [128, 128] | [128, 128] | [128, 128] | [256, 256] | [128, 128] |
Activation Function | ReLU | ReLU | ReLU | ReLU | ReLU |
Learning Rate | 0.0001 | 0.0001 | 0.001 | 0.0001 | 0.0001 |
Training Batch Size | 500 | 2000 | 2000 | 1000 | 500 |
SGD Minibatch Size | 128 | 32 | 32 | 128 | 128 |
Number of SGD Iterations | 10 | 5 | 10 | 10 | 10 |
Number of Training Iterations | 500 | 500 | 800 | 1500 | 500 |
B.4 Evaluation Results of Heuristic Baseline Variants
Desired Inventory | Constant | Variable | Larger | Seasonal | Normal |
-208.00 (0.00) | -435.69 (49.15) | -234.28 (102.81) | -207.75 (34.67) | -150.67 (101.80) | |
-252.00 (0.00) | -479.69 (49.15) | -310.74 (109.16) | -229.08 (34.66) | -226.31 (103.32) | |
-296.00 (0.00) | -523.69 (49.15) | -392.21 (111.79) | -274.29 (40.75) | -322.44 (99.59) | |
-364.00 (0.00) | -390.17 (44.24) | -393.31 (79.05) | -525.84 (47.85) | -283.39 (61.83) | |
-120.00 (0.00) | -395.68 (41.86) | -470.55 (76.73) | -524.26 (64.68) | -351.23 (90.08) | |
-360.00 (0.00) | -412.41 (41.76) | -265.07 (99.67) | -421.90 (55.18) | -232.20 (75.45) | |
-252.00 (0.00) | -382.77 (48.50) | -489.75 (110.96) | -610.03 (94.43) | -177.54 (70.87) | |
-361.00 (0.00) | -397.22 (50.02) | -325.81 (98.39) | -479.07 (69.47) | -218.98 (73.26) |
@@ B.4 Evaluation Results of Heuristic Baseline Variants
Appendix C Case Studies
@@
C.1 Variable Demand Scenario
@@
C.2 Seasonal Demand Scenario